shift summary, evening, Aug. 7, 1999

Jim Hirschauer (jhirsch@cuhep.Colorado.EDU)
Sat, 7 Aug 1999 23:02:08 -0600 (MDT)


Shift Summary for evening shift, 7/Aug/99

On Shift: Theo Alexopoulos, Jim Hirschauer

Hours of beam delivered at all intensities: 8.0
Hours of beam delivered at nominal intensity: 8.0
Hours of E832 data (special runs not included): 7.75
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: .25 hrs - starting new run

========================================================================
Runs ended this shift: 14137
Run 14137 - 24.2M events - BEAM_832
Run 14138 - in progress - BEAM_832

Tapes brought to FCC this shift: QKC221 - QKC226
QKS221 - QKS226
QKK221 - QKK226
QKB081, QKB082

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:

Smooth E832 datataking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14127 15:30 Cryo problems w/ no estimate of time.

295 16:00 Beam returns. The fatted calf is slaughtered.

16:40 Got few spills w/ BOA > 1.5, called MCR. Things
are fine autotune is running.

399 18:25 ktevana errors -> kfrend

411 18:40 ktevana errors -> kfrend (spills 411-413)

21:45 run 14127 ended, tapes full

14128 1-3 22:00 Run 14128 started.
ktevana errors -> kfrend
Noticed missing quadrant on the HCC Illum plot.
Reloaded HCC & Etot and kfrended, everything is
fine. (first 3 spills effected)

22:30 tapes QKC221 - QKC226 vaulted
QKS221 - QKS226
QKK221 - QKK226
QKB081, QKB082

23:00 SEM ~ 7.2E12, called MCR. SEM calms down.

47 23:05 ktevana errors -> kfrend

23:30 Dzero Fiber Tracker Upgrade manual found in
control room. A report will be submitted to the
Dzero group early next week at latest.
========================================================================
Follow up Notes on Previous Shift

o Extraneous assault rifles were taken to FCC, except for the one with a
wood grain handle which Theo insisted was neither extraneous nor an
assault rifle.

========================================================================
Items to Follow up on Next Shift

o Do a muon run & polarity flip at next tape change or beam
disappearance, be it owl or day shift.

o Do a laser scan at beam disappearance > 1.5 hrs.